If angle a is complementary to angle b and the measure of angle b is 36 degrees less than angle a what is the measure of angle a and angle b
Answer by Timnewman(323)

Hi dear,
a+b=90---(1)
b=a-36----(2)
put 2 in 1
a+a-36=90
2a-36=90
2a=90+36
2a=126
a=63
put a=63 in equ 2
b=63-36
b=27
The measure of anle a is 63deg
and the measure of angle b is 27deg.
HOPE THIS HELPS?